An auto repair shop takes an average of 45 minutes to complete a repair job. The owner of the shop has determined the repair time has a standard deviation of 8 minutes. A customer comes into the shop and states that she will return in 35 minutes -- at which time she must have the car ready to drive off. Using the normal probability distribution, what is the probability the car will be repaired in 35 minutes or less?
Answer by stanbon(29603)

An auto repair shop takes an average of 45 minutes to complete a repair job. The owner of the shop has determined the repair time has a standard deviation of 8 minutes. A customer comes into the shop and states that she will return in 35 minutes -- at which time she must have the car ready to drive off. Using the normal probability distribution, what is the probability the car will be repaired in 35 minutes or less?
----------------
Find the z-score of 35.
z(35) = (35-45)/8 = -1.25
P(time < 35) = P(z<-1.25) = 0.1056
====================================
Cheers,
Stan H.